The population of a particular country consists of three ethnic groups. Each individual belongs to one of the four major blood groups. The accompanying joint probability table gives the proportions of individuals in the various ethnic group blood group combinations. Blood Group 0 AB 1 0.082 0.102 0.009 0.004 Fthnic Group 0.139 0.141 0.018 0.005 3 0.215 0.205 0.060 0.020 Suppose that an individual is randomly selected from the population, and define events by A type A selected, B - type B selected?, and C - ethnic group 3 selected. (a) Calculate p(A), P(C), and PA n C). (Enter your answers to three decimal places.) P(A) P(C) (b) Calculate both P(A | C) and P(C | A). (Round your answers to three decimal places) RAIC)- Explain in context what each of these probabilitles represents. Select all that apply.) If a person has type A blood, the probability that he is from ethnic group 3 is given by P(C IA) If a person has type B blood, the probability that he is from ethnic group 3 is given by P(C I A) If we know that the individual cane from ethnic group 3, the probability that he has type A is given by P(C I A) If a person has type B blood, the probability that he is from ethnic group 3 is given by P(A | c). If we know that the individual came from ethnic group 3, the probability that he has type A is given by P(A | C). If a person has type A blood, the probability that he is from ethnic group 3 is given by P(A | C). (c) It the selected individual does not have type B blood, what is the probability that he or she is from ethnic group 1? (Round your answer to three decimal places.)
Answer #1

Solution : ( a )

P(A) 0.102 0.141 0.205 0.448

Р(С.- 0.215 + 0.25 + 0,060 + 0 .020 0.500

P(A n C) = 0.205

=========================================================

Solution : ( b )

If we know that the individual came from ethnic group 3, the probability that he has type A is given by P(A C)0.205 0.410 P(A IC) C0.500

If a person has type A blood, the probability that he is from ethnic group 3 is given by P(C A)

P(AnC) 0.205 P(A) 0.448 0.45758 ︵ะ 0.458

=========================================================

Solution : ( c )

Let E= ethnic group 1 and B not blood type B

Since P(En Bc)0.082 0.102 0.004 0.188

and P(B)= 1-PlB)= 1-(0.009+0.018+ 0.060)= 1-0.087= 0.913

P(EnB) 0.188 P Be0.913 = 0.20591 ะ 0.206